FREQUENCY-COMB REFERENCED, SUB-DOPPLER SPECTROSCOPY OF HOT BANDS OF ACETY-LENE IN THE REGION OF THE ν_{1} + ν_{3} COMBINATION BAND

\underline{CHRISTOPHER P. MCRAVEN}, Chemistry Department, Brookhaven National Laboratory, Upton, New York 11973; MATTHEW J. CICH, SALVATORE M. CAIOLA, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; TREVOR J. SEARS^{a}, Chemistry Department, Brookhaven National Laboratory, Upton, New York 11973; GREGORY E. HALL, Chemistry Department, Brookhaven National Laboratory, Upton, New York 11973.

To take full advantage of recent improvements in the accuracy obtained in collision-induced absorption line shape measurements of the acetylene ν_{1} + ν_{3} combination band, independent and accurate knowledge of the positions of weaker, overlapping hot band transitions is required. We have used cavity-enhanced saturation-absorption spectroscopy together with a frequency-comb referenced, extended cavity diode laser to determine Doppler-free transition frequencies for the ν_{4} and ν_{5} hot band transitions in this region . A laser intensity dependent asymmetry in the sub-Doppler saturation line shape is observed, and will be discussed in terms of the intra-cavity beam geometry. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.

aalso: Department of Chemistry, Stony Brook University, Stony Brook, New York 11794